Larrys Lawn Service needs to purchase a new lawnmower costing $8,476 to replace an old lawnmower that cannot be repaired. The new lawnmower is expected to have a useful life of 4 years, with no salvage value at the end of that period Click here to view the factor table. (a) If Larrys required rate of return is 11%, what level of annual cash savings must the lawnmower generate to be considered an acceptable investment under the net present value method? (For calculation purposes, use 4 decimal places as displayed in the factor table provided and round final answer to O decimal place, e.g 58,971.) Annual cash savings should be (b) If Larrys required rate of return is 14%, what level of annual cash savings must the lawn mower generate to be considered an acceptable investment under the net present value method? (For calculation purposes, use 4 decimal places as displayed in the factor table provided and round final answer to 0 decimal place, e.g 58,971.) Annual cash savings should be
Answer #1

For the lawnmower to be acceptable using the net present value method, the net present value must be greater than or equal to zero. Let X equal the present value of the annual cash saving:

a) (X × PVA(6,11%)) - $8476 = 0

=X × [{1-(1+0.11)-4}÷0.11] = 8476

= X × 3.10244 = 8476

= X = 2732

b) (X × PVA(4,14%)) = 8476

= X × 2.193712 = 8476

X = $2909

For any query please comment and DO GIVE POSITIVE RATING